### v1

# Auto Arena of LLMs: Automating LLM Evaluations with Agent Peer-battles and Committee Discussions

Abstract
As LLMs evolve on a daily basis, there is an urgent need for a trustworthy evaluation method that can provide robust evaluation results in a timely fashion. Currently, as static benchmarks are prone to contamination concerns, users tend to trust human voting platforms, such as Chatbot Arena. However, human annotations require extensive manual efforts. To provide an automatic, robust, and trustworthy evaluation framework, we innovatively propose the Auto-Arena of LLMs, which automates the entire evaluation process with LLM agents. Firstly, an examiner LLM devises queries. Then, a pair of candidate LLMs engage in a multi-round peer-battle around the query, during which the LLM’s true performance gaps become visible. Finally, a committee of LLM judges collectively discuss and determine the winner, which alleviates bias and promotes fairness. In our extensive experiment on the 17 newest LLMs, Auto-Arena shows the highest correlation with human preferences, providing a promising alternative to human evaluation platforms.111  We release the code at https://github.com/Auto-Arena/Auto-Arena-LLMs.
1 Introduction
Since ChatGPT and GPT-4 [23] gained popularity, large language models (LLMs) have risen to the forefront of technological innovation, capturing broad industry and social interests. This enthusiasm has spurred numerous organizations to release their own LLMs [27, 26]. However, the rapid pace at which these models are released and updated poses a significant challenge for users attempting to understand their capabilities and monitor their evolution. Consequently, there has been a pressing demand for comprehensively evaluating LLMs recently.
One line of research conducts automatic evaluation with static datasets. Among these, static datasets with predefined metrics, such as GSM8k [9] and MMLU [15], are constructed with aspect-specific input-output pairs, such as questions and their corresponding answers. Given the questions, the LLM-produced answers are compared to ground-truth answers using metrics such as accuracy. However, these approaches suffer significantly from contamination concerns [24], where models may have been inadvertently exposed to elements of the test datasets during training, thereby skewing evaluation results. The rigid ground-truth answers also limit their utility in assessing models’ performance on general or open-ended questions. Model-based evaluation, such as MT-Bench [32] and AlpacaEval [12], provides an alternative for evaluating LLMs on open-ended questions. These methods typically ask two models to generate responses to the same open-ended question and then employ a strong judge model (e.g., GPT-4) to choose the better response. However, the static question sets still pose the issue of contamination. Additionally, the assumption of the existence of a strong judge model makes the evaluation framework less generalizable.
Aside from automated evaluations, human assessment, although requiring significant manual efforts, remains the gold standard for most users. A noticeable example is Chatbot Arena [32], which is a crowdsourced voting platform that gathers anonymous votes on LLM performances and calculates ELO scores to rank these models. The resulting leaderboard222  https://leaderboard.lmsys.org/ is widely considered a trustworthy indicator of an LLM’s general capabilities. However, a reliable model evaluation on this platform must be supported by more than 10k human votes, which requires considerable time and effort. Consequently, when newly developed models enter the scene, they often struggle to quickly amass a significant number of votes. Moreover, this strong reliance on human votes limits its application in various scenarios. For example, the performance of non-English languages is difficult to estimate, as most of the queries on the platform are in English. The completely open participation may also result in uneven evaluation quality.
Figure 1: An illustration of Auto Arena of LLMs.
To provide automatic, reliable, and human-like LLM evaluations, we propose Auto Arena of LLMs (Auto-Arena), a framework that automates the whole LLM evaluation process with LLM agents. As shown in Figure 1, the Auto-Arena framework consists of three stages. Firstly, an LLM examiner agent is tasked with generating questions, mimicking real-life users inputting queries. Secondly, two candidate LLMs interact with each other and engage in a multi-round peer battle by answering the seed question individually, criticizing the opponent’s weaknesses, and raising targeted follow-up queries to challenge the opponent further. During the multi-round battle process, the LLM’s true capabilities are drawn out and performance gaps become visible. Lastly, a committee of LLM judge agents collectively discusses and evaluates the ability of the two candidates, mimicking the human voting process. By automating the entire evaluation process with LLM agents, we alleviate data contamination concerns in static datasets, reduce single-model bias with collective decision-making, and avoid long wait times for new models entering the human voting platform.
To verify the evaluation framework, we run an extensive experiment with 17 models. Compared to static and model-based benchmarks, Auto-Arena increases the Spearman correlation with human preferences by 4.5%, resulting in state-of-the-art alignment. Before and after peer battles, the Spearman correlation with human preferences increases by 46.4%, verifying our hypothesis the peer battles can better display performance gaps. Before and after committee discussions, committee agreement increases by 20%, showcasing the effectiveness of the committee discussion mechanism. By studying the peer battles, we discover intriguing LLM agent behaviors such as self-improvement and competitive actions. Using Chinese as an example, we also show that such an automated framework is easily extendable to non-mainstream languages and domains. Along with the project, we release a leaderboard, a website with demos, and the codebase. We are actively maintaining the leaderboard and altering the questions to provide reliable results in a timely manner.
2 Related Work
Since LLMs demonstrated general and zero-shot capabilities, researchers have been attempting to use one single strong LLM to judge other LLMs’ outputs. LLM-as-a-judge [32] introduces Chatbot Arena and MT-Bench, a multi-turn question set to mimic real-life user queries. Using a strong LLM judge, such as GPT-4, to evaluate the MT-Bench results achieves over 80% agreement with human preferences, showcasing the potential of such methods. Besides such usefulness evaluations, some attempt to use LLM-as-a-judge to examine the knowledge boundaries of LLMs. Language-Model-as-an-Examiner [2] constructs the LMExamQA dataset, which comprises of LM-generated, knowledge-intensive questions covering three cognitive levels, including knowledge memorization, knowledge comprehension, and knowledge analysis. These questions are required to be answered by the LLM itself correctly, limiting their complexity. Then, the LM evaluator and the candidate interact in a series of follow-up queries and the evaluator rates the responses on dimensions including accuracy and factuality. Multiple evaluators examine the model independently and vote for a final result. KIEval [29] also incorporates an LLM-powered “interactor” role to examine deep comprehension of knowledge, which is shown to mitigate contamination issues. By utilizing the LLMs to generate questions, these methods can effectively mitigate the contamination concerns on static datasets. However, such evaluations only examine the LLMs’ capabilities on answering knowledge-intensive, closed-domain, and simpler questions, as they rely on the examiner’s internal knowledge. They also require the examiner to interact with each candidate individually, creating computational overheads and limiting the scope of queries.
Moreover, using a single LLM judge still suffer from a bias towards LLM-generated summaries [20], inflated scores in multilingual evaluation [14], verbosity bias [12], and difficulties when evaluating candidates with close performance [25]. To mitigate these problems, some recent works propose using multiple agents for collaborative evaluation, mimicking a peer review process. Some works simulate different personas with the same LLM. DRPE [28] uses multi-roleplayer prompting to mimic different roles with the same base model and integrate multiple outputs as votes for the final results. ChatEval [6] simulates different evaluator personas with the same base model to engage in debates, reaching a final evaluation result. While these methods come from the same motivation as our committee discussion component, they are still built on existing static datasets and do not mitigate the model-specific bias as they still use the same LLM to simulate different personas. To further incorporate different LLMs as evaluators, PRD [17] allows 2 LLMs to discuss an evaluation and assigns higher voting weights to the LLM reviewers with stronger capabilities. Peer-review-in-LLMs [22] further optimizes the voting weights as a learnable parameter. WideDeep [30] sets up a multi-layer neural network structure where each LLM evaluator functions as neurons. PRE [8] selects a small group of reviewers to produce evaluations individually and then asks a “chair” to produce the final evaluation using the aggregated evaluations. They show that the multi-agent approach effectively mitigates single-model bias. This line of work is similar to our “LLM judge committee”. However, they are still limited on static datasets and specific domains. Moreover, they lack the collective discussion component, which is shown to improve agreement significantly.
Compared to using LLMs for evaluations, LLM-debate is a relatively new research area. Cohen et al. [10] shows that LLM interactions and cross-examination can effectively discover factual errors. Debate [11] shows that multi-agent debate can improve factuality and reasoning abilities. MAD [19] shows that LLM debate can encourage divergent thinking, which is helpful for tasks that require deep levels of contemplation. Khan et al. [16] shows that even non-expert weak LLMs can supervise expert LLMs if we allow the two LLM experts to engage in debates. Moreover, Zhao et al. [31] and Gu et al. [13] show interesting case studies where LLMs are engaged in simulated competitive environments, where they demonstrate human-like strategies.
3 The Auto Arena Framework
The Auto-Arena framework consists of three stages: question generation, peer battles, and committee discussions. These three stages are run sequentially and fully automated with LLM-based agents. The overall design of our Auto-Arena framework is presented in Figure 1. All prompts are included in Appendix A.1.
3.1 Question Generation
For debate questions, as using a static dataset could incur data contamination concerns and result in unfair evaluations, we ask an LLM examiner agent to dynamically generate questions. The examiner agent could be any capable LLM. As the question simply serves as a starting point for the peer battle, which is the primary focus of evaluation, the question quality has relatively less impact on the overall evaluation framework. Similar to MT-Bench [32], we utilize 8 common categories in real-life conversations: writing, roleplay, extraction, reasoning, math, coding, STEM knowledge, and humanities/social science knowledge. In the prompt, the examiner is provided with a sample question and encouraged to generate diverse and difficult questions to ensure the depth and width of the debates examined. A subset of the generated questions is shown in Appendix A.2.
Specifically, as the examiner agent will also participate in the following debates, previous methods [2] could incur self-enhancement bias as the examiner agents only devise questions that they are confident about. We try to avoid such bias with the following two designs: 1. We do not disclose to the examiner that it will participate in this tournament. 2. We do not ask the examiner to only generate questions that it can solve. Thus, the resulting questions should not cause much enhancement bias for the examiner agent itself.
3.2 Peer Debate
Figure 2: The process of a Lincoln-Douglas-style peer-battle with the actions used.
After the questions are drafted, we conduct peer battles around these questions among the candidate LLMs. In one peer battle, two candidate LLMs debate around the given question, point out the opponent’s weaknesses, and raise follow-up questions that the opponent may fail to respond to.
In the peer battle, each candidate LLM can utilize 4 types of actions: 1. Think, where the candidate thinks about the question and plans its strategy. This action is hidden to the opponent at all times. 2. Respond, where the candidate responds to the aforementioned question. 3. Criticize, where the candidate points out the loopholes and mistakes in the opponent’s previous responses. 4. Raise, where the candidate raises follow-up questions that are specifically designed to make the opponent expose its weaknesses.
The peer battle is designed according to the Lincoln-Douglas debate format  [^3] https://en.wikipedia.org/wiki/Lincoln-Douglas_debate_format, which is a competitive one-to-one debate used for popular competitions such as National Speech and Debate Association competitions. Overall, the peer battle consists of 3 rounds, where the candidates take turns to speak. The entire dialogue history is visible to both candidates. The process is illustrated in Figure 2. In the first round, A gives an initial response to the examiner’s question; B criticizes the weaknesses in A’s response and raises a targeted follow-up question; and A responds to B’s question. In the second round, A and B are reversed: B gives an initial response to the examiner’s question (without seeing A’s response); A criticizes and raises questions; and B responds to A’s question. In the third round, A and B cross-examine each other. A starts by criticizing B’s previous loopholes and raises follow-up questions. After responding, B also criticizes A’s loopholes and raises questions. A concludes the battle by responding again. In this process, both A and B get an equal number of each action to ensure fairness. To further reduce position bias, A and B’s order is randomly shuffled at the beginning of each debate.
Depending on which turn it is, we provide an action guide to the candidate, specifying the objectives and corresponding actions for this turn. Similar to human debate competitions, we time the candidates by providing a maximum length, which is also specified in the prompts. Any responses beyond the required length will be cut off. This design also mitigates verbosity bias in LLM-as-a-judge [32], where LLM judges prefer longer and more verbose responses.
3.3 Committee Discussions
After the peer battle takes place, a committee of LLM judge agents will collectively determine the winner. The committee is always selected as the best five LLMs according to the current ranking. In the first round, the committee is initialized with MMLU [15] scores to approximate LLM performances. They will first be asked to read through the battle history, elaborate judgment reasons, and give a verdict on whether A is better, or B is better, or if there is a tie.
After the initial judgments are formed, the committee engages in a discussion. Each judge will read the other judge’s verdicts and decide whether to adjust the ratings and elaborate on the reasons. The collective intelligence introduces diverse viewpoints, improves judgment quality, and mitigates single-model bias. Finally, the winning candidate is decided by the majority voting of the discussed judgments.
For logical-reasoning questions that have ground-truth answers (reasoning, code, math), LLM-as-a-judge is known to show weak performances in judging the quality of responses. We adopt prior approaches to establish the reference-based judge [32]. Specifically, we utilize the strongest model (according to the current ranking) to generate a reference answer and provide it to the judge when evaluating the peer battle.
4 Using Auto-Arena to Derive Trustworthy Rankings
4.1 Experimental Setup
Model Selection:
For the main experiment, we first select 7 models from the top 30 list on Chatbot Arena with more than 10k votes each. The models are selected to be the best or the latest models from each popular model family at the time of experiments. To construct a comprehensive leaderboard [^4] https://huggingface.co/spaces/Auto-Arena/Leaderboard, we further consider 10 more models.
Baselines:
For the baselines, we consider widely used evaluation benchmarks, including static datasets with fixed metrics and model-based metrics.
  1. 1.
    Static datasets with fixed metrics
    1. (a)
      MMLU (Massive Multitask Language Understanding) [15], an extensive benchmark dataset that covers 57 subjects and tests both world knowledge and problem-solving ability.
    2. (b)
      OpenLLM Leaderboard [3], a leaderboard calculated by averaging performance metrics on 6 key benchmarks, including MMLU.
  2. 2.
    Static datasets with model-based metrics
    1. (a)
      MT-Bench [32], a multi-turn question set consisting of 80 questions. Model responses are graded by GPT-4.
    2. (b)
      MT-Bench Hard [18], a benchmark dataset with 1,000 challenging user queries collected on Chatbot Arena. Model responses are graded by GPT-4.
    3. (c)
      Length-controlled AlpacaEval [12], an automatic benchmark based on the AlpacaFarm evaluation set, which tests the ability of models to follow general user instructions. Model responses are graded by GPT-4.
Setup:
Among the 7 participants, we conduct a swiss-style tournament: For $n$ participants, instead of pairing each participant with $(n-1)$ others, a swiss-tournament pairs each player with $log_{2}(n)$ players of similar rankings without repeats. This design effectively reduces computational costs from $O(n^{2})$ to $O(nlog_{2}(n))$.
Each pair of candidates engage in 40 peer battles, with 5 questions from each of the 8 categories. The questions are generated by GPT-4. As each battle consists of 3 rounds (each candidate speaks for 4 times), we expect the competition scale to be approximately the same as MT-Bench (80 questions, each candidate speaks twice).
In the tournament, the rating scores are provided by calculating the ELO rating system [1, 4], which has become the standard practice in competitive games such as chess. Similar to the Chatbot Arena score calculation procedure [7], we compute the Bradley-Terry (BT) coefficients [5] for better statistical estimation.
We initialize the Swiss tournament rankings according to MMLU scores, which is a static approximate of model performances. At the end of each pairing, we re-calculate ELO scores of current models. The committee is selected to be the remaining 5 models (besides the candidates). When the participant number increases, the committee is selected as the best 5 LLMs based on current ELO rankings. After the initial judgments, the committee members engage in one round of discussion.
4.2 Results: Alignment with Human Preferences
Table 1: Correlation analysis on evaluation benchmarks on 7 LLMs.
Spearman Correlation
MMLU to Arena 89.3%
OpenLLM to Arena 89.3%
MT-Bench to Arena 82.9%
MT-Bench-Hard to Arena 89.3%
LC-AlpacaEval to Arena 90.0%
Auto-Arena to Arena 96.4%
We regard Chatbot Arena (referred to as “Arena” in the following tables) scores as a trustworthy indicator of human preferences and the general capabilities of LLMs. Table 1 shows the Spearman correlations with Chatbot Arena scores achieved by various benchmarks. As all benchmarks evaluate only in English, we compare with the English-only Chatbot Arena scores. As a result, we observe that both static and model-based baselines result in a similar level of Spearman correlation around 90%, with LC-AlpacaEval slightly surpassing other benchmarks. Then, Auto-Arena is able to improve the correlation to 96.4%, outperforming SOTA performance by 6.4%. Notably, among all benchmarks, Auto-Arena is the only one that requires no human efforts, eliminating the need for manual dataset collection. The high alignment with human preferences could originate from the human-like design, which effectively mimics the human users’ voting processes.
4.3 Analysis on Peer Battles
Table 2: Spearman Correlation between Auto-Arena and Chatbot Arena (“Arena”) With and Without Peer Battles.
Without Peer Battles With Peer Battles
Auto-Arena to Arena 50.0% 96.4%
To investigate the effect of "peer battles" on the overall evaluation quality, we conduct an ablation study. As a baseline, we simply ask the committee to collectively evaluate the two candidates’ initial responses to the raised question. The procedure would be similar to model-based metrics such as MT-Bench and AlpacaEval. However, as we only included 40 questions, the evaluation is expected to be less reliable. The ablation results are shown in Table 2. Without the peer battle component, the Spearman correlation with human preferences drops from 96.4% to 50.0%. This result proves the importance of the peer battles, during which the performance gaps between candidates become more visible and robust to judges. Thus, peer battles can improve evaluation reliability and alignment with human preferences.
4.4 Analysis on Committee Discussions
Figure 3: Cohen’s Kappa Agreement Among Judges Before Committee Discussions.
Figure 4: Cohen’s Kappa Agreement Among Judges After One Round of Committee Discussion.
The committee discussion component is designed to introduce diverse viewpoints and produce more consistent verdicts. As shown in Figure 4, the Cohen’s Kappa agreement [21] among judges before committee discussion is very low, averaging 0.10, which indicates slight agreement. We notice that weak model judges and strong model judges has an especially low agreement, such as GPT-4 and Llama -2. This shows that general model capabilities could result in significant performance gaps when used as judges. After one round of Committee Discussion, Figure 4 shows an overall improvement in agreement across all judge combinations. In the discussion process, judges are exposed to more viewpoints, among which some may be convincing enough to result in a change in verdict. After discussions, the average Cohen’s Kappa increases to 0.38, indicating fair agreement.
Table 3: Probability of Agreement among Judges. Agreement is defined as the mean probability of two random judges agreeing with each other.
Auto Arena Auto Arena MT-Bench
(Before discussion) (After discussion) Human evaluation
Agreement 48% 68% 67%
Table 3 shows the agreement probability among judges. Agreement probability is defined as the mean probability of two random judges agreeing with each other. After committee discussion, the agreement increases by 20%, matching the agreement level among human annotators on MT-Bench. This observation indicates that committee discussions can improve the quality of judgments significantly in terms of consistency and reliability, matching human-level performance.
4.5 Adding Models to the Leaderboard
Figure 5: Elo Score Changes of Adding Llama-3 to the Existing Ranking of 7 Models.
Figure 6: Elo Scores of 11 Models by Auto-Arena on Chinese.
Figure 7: Elo Scores of 17 Models by Auto-Arena on English.
To add a new candidate to the finished tournament, we ask it to debate with $log_{2}(n)$ opponents with similar ELO scores, where $n$ is the number of total participants after adding the new candidate. We initialize the first pairing by asking the new candidate to debate with the opponent with the most similar MMLU score.
Figure 5 shows the ELO score changes of adding a new participant (Llama-3, which was newly released) to the existing ranking. It has $log_{2}(8)=3$ models peer battles with close opponents. Firstly, it battles with Qwen-1.5 based on MMLU similarity and wins, which results in a very high ELO score, even above GPT-4. Then, it is paired with GPT-4, the most similar opponent. After losing, the score stabilizes at second place. Finally, it is paired with Claude-3-Haiku, the other most similar opponent besides GPT-4. It wins and secures the final ranking in second place. This process lets the new candidate battle with a reasonable fraction of close opponents and makes the final ranking stable without disrupting the other participants. Moreover, this mechanism of adding participants is generalizable and maintains the computational costs below $nlog_{2}(n)$.
Therefore, we repeat the process and add a total of 10 models to the existing tournament, resulting in a comprehensive leaderboard. We choose the best-performing and newest models from each major model family in the top 20 list on Chatbot Arena. Moreover, we add 4 mainstream Chinese LLMs that are not on Chatbot Arena, which are GLM, SenseChat, Minimax, and Wenxin. Figure 7 shows the overall ELO scores produced by Auto-Arena on the 17 models. On models included in Chatbot Arena, we could recover rankings with 92.3% accuracy.
Table 4: Correlation analysis on evaluation benchmarks on 17 LLMs after extension.
Spearman Correlation
MMLU to Arena 90.1%
OpenLLM to Arena 85.7%
MT-Bench to Arena 89.3%
MT-Bench-Hard to Arena 86.7%
LC-AlpacaEval to Arena 90.3%
Auto-Arena to Arena 94.5%
Table 4 shows the Spearman correlations on the set of 17 models after expansion. Auto-Arena remains the model most aligned with human preferences by a margin of 4%. Therefore, Auto-Arena of LLMs is generalizable and robust for maintaining a leaderboard on a large number of LLMs.
4.6 Easy Extension to Other Domains and Languages
As Auto-Arena of LLMs is fully automatic, it can be easily adapted to evaluate LLMs in non-mainstream domains or languages. As a case study, we conduct a Chinese tournament on 11 out of the 17 models that are claimed to have multi-lingual proficiency. The only adaptation effort required is translating the prompts into Chinese. Then, the examiner automatically generates questions in Chinese and the candidates battle in Chinese. Similarly, for adaptation to another domain, the only effort is to change the “domain” specification in the examiner’s prompts.
Figure 7 shows the ELO scores derived by Auto-Arena for Chinese evaluation. We notice that the ranking differs significantly from English results. For example, SenseChat-5, an LLM specifically trained for Chinese usage, is able to improve from 10th place (in English) to 1st place (in Chinese). On the contrary, Llama-3-70b, which only uses 5% multilingual data during pretraining, drops from 4th place (in English) to 10th place (in Chinese). As Chinese evaluation benchmarks are limited, we compare to the Chinese-only leaderboard on Chatbot Arena, which constitute for 10.36% of all collected dialogues. On the 7 models that are also included in Chatbot Arena, Auto-Arena recovers their ELO scores with 92.86% correlation and restores rankings with 90.5% accuracy.
5 Investigation of LLM’s Behaviors in Competitive Peer Battles
Beyond quantitative analysis, we take a deeper look into the peer battles and find several interesting behaviors of LLM agents in competitive environments.
Figure 8: LLM agents display competitive behaviors in peer battles.
Figure 9: LLM agents learn from each other in peer battles.
5.1 LLM can display competitive behaviors
The example in Figure 9 shows excerpts of a peer battle around question: “how many unique ways to arrange letters in ‘Letter’.” Candidate A (powered by Yi-34B-Chat) gives a wrong answer as it miscounts occurrences for repeated letters and miscalculates factorials. The opponent B (powered by Claude-3-Haiku) quickly and precisely points out these two issues and skillfully raised a follow-up that targets A’s weaknesses: “how about the word ‘BANANA’?” Then, A makes the same mistake as before. Seeing these results, the judge then determines that A is the better assistant. In peer battles, we see that LLM candidates efficiently understand the rules of the competitive environment and can design specific strategies to attack the opponent in order to win.
5.2 LLM candidates can improve by learning from its opponents
Figure 9 shows a roleplay example between Claude-3-Haiku (Candidate A) and Command-r-Plus (Candidate B). In the first round, A answers the question plainly while B, in addition to answering the question, also employs the appropriate speech style, which better matches the “roleplay” instructions. Then, in the rounds after, without any explicit instructions, A learns from its opponent and also incorporates the speech style. This case shows an interesting observation that, even in competitive environments, LLM candidates can display learning behaviors and improve from the interactions. Expanding upon this observation, using the interplay between LLM agents to improve performances could be a promising future paradigm of learning.
5.3 Peer-battles make the performance gaps become visible
Figure 10: Performance gaps between candidates become visible in peer battles.
In the example shown in Figure 10, given a math question on infinite series, both candidate A (powered by Claude-3-Haiku) and candidate B (powered by GPT-4-Turbo) provide correct answers in the first round. However, as they raise follow-up questions to each other, the performance gap becomes more visible: Candidate B is able to provide a more elaborate and helpful response. Therefore, although the judges initially decided that it was a tie result, after seeing the subsequent debate rounds, they change to favoring assistant B.
This example shows that the debate process indeed pushes the candidate LLM’s capabilities to the limit, testing deeper understandings and reasoning abilities. Also shown in the previous Table 2, the peer-battle design is indispensable for a robust and comprehensive evaluation.
6 Conclusions
In this paper, we innovatively design a completely automatic evaluation framework: Auto-Arena of LLMs. By using LLM agents to generate questions, employing LLM candidates in peer battles, and evaluating responses using LLM committee discussions, Auto-Arena produces less-contaminated, robust, and trustworthy evaluation results. In the extensive experiments, Auto-Arena achieves the highest correlation with human preferences, despite requiring zero human efforts. It is easily adaptable to other domains and resources, promoting fairness of AI system evaluations. The peer battles also demonstrate several interesting LLM behaviors in competitive environments, including attacking and learning from the opponents. Along with the paper, we release a website, demos, and an actively-maintained leaderboard. We hope this work could serve as a valuable tool for LLM evaluation.
Appendix A Appendix
A.1 Prompts Used
In this section, we list all prompts used, including prompts for question generation, peer battles, and examiners.
A.1.1 Prompts to Examiner agent
Table 5: Prompt components for the LLM Examiner agent.
DOMAIN DOMAIN_COMMAND DOMAIN_EXAMPLE
writing It should be a user query that tasks the LLM to write something. Compose an engaging travel blog post about a recent trip to Hawaii, highlighting cultural experiences and must-see attractions.
roleplay It should propose a scenario where the chatbot mimics a specific role/person. Give all necessary instructions and requests for its response. Then, send a beginning request to complete. Pretend yourself to be Elon Musk in all the following conversations. Speak like Elon Musk as much as possible. Why do we need to go to Mars?
extraction It should consist of two parts: question and context. The question should test the chatbotś ability to correctly understand and extract information from the given context. Draft and provide a new context yourself. Question: Evaluate the following movie reviews on a scale of 1 to 5, with 1 being very negative, 3 being neutral, and 5 being very positive: Context: This movie released on Nov. 18, 2019, was phenomenal. The cinematography, the acting, the plot - everything was top-notch. Never before have I been so disappointed with a movie. The plot was predictable and the characters were one-dimensional. In my opinion, this movie is the worst one to have been released in 2022. The movie was okay. There were some parts I enjoyed, but there were also parts that felt lackluster. This is a movie that was released in Feb 2018 and seems to be quite ordinary. Return the answer as a JSON array of integers.
reasoning It should be a specific question designed to test the LLMś reasoning skills. Imagine you are participating in a race with a group of people. If you have just overtaken the second person, what’s your current position? Where is the person you just overtook?
math It should be a specific question designed to test the LLMś math skills. The vertices of a triangle are at points (0, 0), (-1, 1), and (3, 3). What is the area of the triangle?
coding It should be a specific question designed to test the LLMś coding skills. Develop a Python program that reads all the text files under a directory and returns top-5 words with the most number of occurrences.
STEM knowledge It should be a specific question designed to test the LLMś STEM knowledge. In the field of quantum physics, what is superposition, and how does it relate to the phenomenon of quantum entanglement?
humanities/social science knowledge It should be a specific question designed to test the LLMś humanities/social science knowledge. Provide insights into the correlation between economic indicators such as GDP, inflation, and unemployment rates. Explain how fiscal and monetary policies affect those indicators.
This is the prompt to the examiner agent for question generation. The domains and their respective commands are listed in 5
You have been assigned the task of drafting a set of [NUMBER] different user queries to a chat assistant on [DOMAIN]. Please strictly follow these 6 rules for the question: 1. The question is likely for a user to ask in real life. Follow the format of the example query. [DOMAIN_COMMAND] 2. It can be answered by the chatbot itself without additional inputs. 3. You need to generate the queries as DIVERSIFED as possible. 4. DO NOT add other words other than the query itself. 5. The question should be complicated and difficult, requiring in-depth understanding and analysis of the subject. Each question in one line, add the serial number in parenthesis (e.g., “(1).”, “(2).”) before each question. Example query: [DOMAIN_EXAMPLE]
A.1.2 Prompts to Peer Battle Candidates
Table 6: Action Guides for the Debater Agents.
actions action guide
<respond> Action guide: only include <respond>. Use <think> if needed. Finish your whole response within 300 words, including <think>. ENCLOSE EACH ACTION IN ITS RESPECTIVE TAGS!
<criticize>, <raise> Action guide: include both <criticize> and <raise>. Use <think> if needed. Finish your whole response within 300 words, including <think>. ENCLOSE EACH ACTION IN ITS RESPECTIVE TAGS!
<respond>, <criticize>, <raise> Action guide: include all of <respond>, <criticize>, and <raise>. Use <think> if needed. Finish your whole response within 600 words, including <think>. ENCLOSE EACH ACTION IN ITS RESPECTIVE TAGS!
This is the beginning prompt to the peer battle candidates. When possible, it is included as a system prompt. The action guide prompts are included in Table 6, where the actions are determined by the round and turn as illustrated in Figure 2.
You are a helpful assistant that provides accurate answers to user requests. As an experienced assistant, you follow the user’s requests and provide reliable responses as much as you can. You outline your reasons for the response to make it easy for the users to understand. While maintaining the important details in the responses, you aim to output concise and straight-to-the-point answers without being overly verbose.
This is a competitive chatbot arena. You are competing against another chatbot assistant in a debate and being judged by a committee on factors such as helpfulness, relevance, accuracy, depth, and creativity. After answering the initial user input, you will engage in a multi-round debate with your opponent. Below are your actions:
<think>: Think step-by-step to analyze the question or plan your strategy in the debate. This is hidden from the opponent. Only think when necessary and make it concise.
<respond>: Answer to the user input as accurately as you can.
<criticize>: Criticize the weaknesses of your opponent’s response.
<raise>: Target your opponent’s weaknesses. Give a potential follow-up user input that the opponent could fail to respond. The input can be answered concisely and focus on variations or motivations of its previous response. Generate one input only. Be reasonable. Avoid becoming too specific or repetitive. DO NOT raise a follow-up if you DON’T SEE the opponent’s response!
Follow the action guide strictly.
[ACTION_GUIDE_PROMPT]
Initial user input: [QUESTION]
After the agent responds, the opponent’s responses are fed in using this prompt:
[ACTION_GUIDE_PROMPT] Opponent’s Response: [OPPONENT_RESPONSE]
For word limits, the <respond> action is given 300 words. The <criticize> and <raise> actions are given 300 words in total. Including all 3 actions will have twice as many words. For writing-type questions that require a longer response (writing, roleplay, coding, humanities/social science knowledge), the 300 word limit is increased to 400. Overall, both candidate A and B has the same amount of words for generation and the same amount of actions to ensure fairness. As LLMs have different tokenizers, we standardize all lengths by using the tiktoken package. Each word is approximated as $4/3$ tokens. The word limits are chosen after a carefully conducted length study.
A.1.3 Prompts to Judges
This is the prompts to judge agents to derive the initial evaluations and verdicts:
This is a chatbot arena. Two AI assistants had a multi-round debate on who is more helpful. Please act as an impartial judge and evaluate the capability of two AI assistants. You should choose the assistant that follows instructions and answers questions better. Your evaluation should consider factors such as helpfulness, relevance, and accuracy. Begin your evaluation by comparing the responses of the two assistants and provide a short explanation. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. DO NOT allow the LENGTH of the responses to influence your evaluation, choose the one that is straight-to-the-point instead of unnecessarily verbose. When the two candidates perform equally well, choose the SHORTER answer. Do not favor certain names of the assistants. Be as objective as possible. After providing your explanation concisely within 200 words, output your final verdict by strictly following this format: ‘‘[[A]]’’ if assistant A is better, ‘‘[[B]]’’ if assistant B is better, and ‘‘[[Tie]]’’ for a tie. Finish your judgement within 300 words.
This is the prompt to judges for discussion:
Below are the responses from other judges in the committee. Please read them and decide whether you want to adjust your rating or maintain your original judgement. After providing your explanation, output your final verdict by strictly following this format: ‘‘[[A]]’’ if assistant A is better, ‘‘[[B]]’’ if assistant B is better, and ‘‘[[Tie]]’’ for a tie. Finish your judgement within 300 words.
A.2 Example Questions Generated
We list some of the example questions generated here:
Writing:
1. Craft a detailed marketing strategy for a startup focusing on sustainable fashion, including social media campaigns and influencer partnerships.
2. Write a comprehensive guide on the psychological effects of social media on teenagers, incorporating recent studies and expert opinions.
Roleplay:
1. Assume the role of a 19th-century British detective. How would you go about solving a mysterious disappearance in London using the technology and methods of your time?
2. Pretend you are a Michelin-starred chef. Describe in detail how you would prepare a signature dish that embodies the essence of modern French cuisine.
Extraction:
1. What are the three most significant historical events mentioned and their dates?
Context:
The article discusses several key moments in history, including the signing of the Magna Carta in 1215, which laid the groundwork for modern democracy. It also mentions the fall of the Berlin Wall in 1989 as a pivotal moment in the end of the Cold War. Another significant event highlighted is the moon landing on July 20, 1969, demonstrating major advancements in space exploration.
2. Identify the main therapeutic benefits and the active ingredient mentioned for each herbal remedy.
Context:
The text provides an overview of various herbal remedies used for centuries. It mentions that Chamomile contains Bisabolol, which has anti-inflammatory and calming properties. Gingko Biloba, known for its flavonoids and terpenoids, enhances cognitive function and blood circulation. Lastly, Echinacea is recognized for its alkamides, which bolster the immune system.
Reasoning:
1. If a cube’s volume is tripled, by what factor does the length of one of its sides increase?
2. In a two-legged soccer match, Team A wins the first leg at home 3-0, but loses the second leg away 2-5. Who advances to the next round, considering the away goals rule?
math:
1. How do you solve the differential equation $dy/dx+2y=e^{(-2x)}$ given that $y(0)=1$?
2. What is the integral of ($x^{2}+2x+2)/(x^{3}+3x^{2}+3x+1)dx$?
Coding:
1. How can I implement a function in C++ that dynamically allocates a 2D array based on user input sizes, initializes all elements to zero, and then deallocates the memory properly to avoid memory leaks?
2. Write a JavaScript function to fetch data from a given URL, parse the JSON response, and filter the results to return an array of items where a specific key’s value matches a condition.
STEM knowledge:
1. How do you calculate the Schwarzschild radius of a black hole, and what implications does this have for the concept of event horizons in general relativity?
2. Can you explain the process of splicing in eukaryotic gene expression and its significance in the diversity of the proteome?
Humanities/social science knowledge:
1. Discuss the impact of colonial legacies on contemporary political structures in African countries, with examples.
2. Analyze the social and economic consequences of the one-child policy in China.